OpenStudy (anonymous):

Two balls have the same mass. One ball moving along the positive x-axis and the other ball moving along the positive y-axis undergo perfectly inelastic collision. What can you say about the motion of the balls after the collision?

OpenStudy (irishboy123):

\[m \dot x_1 + m (0) = 2 m \dot x_2 \implies \dot x_2 = \frac{1}{2} \dot x_1\] \[m \dot y_1 + m (0) = 2 m \dot y_2 \implies \dot y_2 = \frac{1}{2} \dot y_1\] 

OpenStudy (irishboy123):



OpenStudy (irishboy123):

symmetry?